In a survey of Jawas, 120 said that they buy and fix all the droids they sell and 230 admitted to stealing droids if the droid is not accompanied by an owner. If one Jawa is selected at random, find the probability that the Jawa has stolen a droid.
Answer #1

Let A = Event that the Jawa buy and fixes all the droids he sells

B = Event that the Jawa steals droids if the droid is not accompanied by an owner

Total no. of subjects surveyed = 230+120 = 350

If one jawa is selected, P(B) = 350

= \frac{230}{350}=0.66

Hence, if one Jawa is selected at random, the probability that the Jawa has stolen a droid is 0.66.